OpenStudy (anonymous):

Help please, medal and fan :) Can someone help out a bit? The following function represents the production cost f(x), in dollars, for x number of units produced by company 1: f(x) = 0.15x^2 - 6x + 400 The following table represents the production cost g(x), in dollars, for x number of units produced by company 2: x | g(x) 50 | 75 60 | 60 70 | 55 80 | 60 90 | 75 The minimum production cost for company _____ is greater. [Put 1 or 2 in the blank space]

OpenStudy (anonymous):

@amistre64 @hartnn can you guys help?

OpenStudy (anonymous):

@ganeshie8 @Hero help please?

ganeshie8 (ganeshie8):

you're given two functions, and asked to tell which one has minimum value... where are u stuck exactly ?

OpenStudy (anonymous):

I'm stuck on how to find the minimum value of the first function, since its not in a table @ganeshie8

ganeshie8 (ganeshie8):

okay, u figured out the minimum value for table already ?

ganeshie8 (ganeshie8):

For the first function, use below : minimum/maximum value for a quadratic function = \(\large f\left(-\dfrac{b}{2a}\right)\)

ganeshie8 (ganeshie8):

find out -b/2a, and evaluate the function at this value

OpenStudy (anonymous):

Ok, so the minimum value for the first function would be 20? And the minimum value for the table is 55, correct? @ganeshie8

ganeshie8 (ganeshie8):

f(x) = 0.15x^2 - 6x + 400 minimum value occurs at x = 20, to find the minimum value, plugin x=20 above and evaluate

ganeshie8 (ganeshie8):

f(x) = 0.15x^2 - 6x + 400 f(20) = ?

OpenStudy (anonymous):

I got..340? I'm not sure if that sounds correct.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Excellent !! 340 is the minimum value for first function

OpenStudy (anonymous):

So, Company 1 has a greater minimum? @ganeshie8

OpenStudy (anonymous):

Also, can you help with one more question?

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

sure, close this ad post a new ... il be back in 10 mnts...

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

@ganeshie8